OpenStudy (anonymous):

plz prove log(1+x)/(1-x) an odd function

OpenStudy (anonymous):

I dont think it is :|

OpenStudy (anonymous):

odd function is if -f(x)=f(-x) this if f(-x)=log(1-x)/(1+x)=-f(x)

OpenStudy (anonymous):

so it is odd function isn't it

OpenStudy (anonymous):

\[f(x)=\ln(1+x)-\ln(1-x)\]

OpenStudy (anonymous):

you are saying that (1-x)/ (1+x) = - [ (1+x) / (1-x) ]

OpenStudy (anonymous):

replace x by -x to get \[f(-x)=\ln(1-x)-\ln(1+x)=-(\ln(1+x)-\ln(1-x))\] so odd